40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1985     Entry time: Thu Jul 23 16:35:05 2015
Author: Arjun 
Type: Misc 
Category: PD noise 
Subject: PD noise update 

As mentioned in a previous eLog, due the inconsistency observed, I decided to measure the Free running RIN and supressed RIN again. I present the results below. A few things are bothering me though:

1)The peak at 25kHz has been supressed! How? UGF is around 10kHz!

2) The supression is more almost 2 orders(40dB) at 100Hz. Expected supression about 30dB.

 The splicing has resulted a sudden jump at 100Hz, but I think thats nothing to be worried about, as expected we get an a little more than 1 order of suppression, in the bandwidth of interest from 3Hz-10kHz. Also, I think I figured out what went wrong last time. I had the spectrum analyzer in the DC coupled mode when I was measuring the free running RIN of the laser, where as for the suppressed case I was using AC coupled mode of spectrum analyzer, I think that could possibly be the reason for the inconsistency observed in my previous measurement.

 

Attachment 1: Intensity_supp.fig  69 kB
Attachment 2: Intensity_supp.pdf  17 kB  | Hide | Hide all
Intensity_supp.pdf
ELOG V3.1.3-